OpenStudy (thcsbabygirl1224):

What is 120 degrees in radians?

OpenStudy (thcsbabygirl1224):

@AloneS. so i would divide this by 180 ?

OpenStudy (aveline):

It's about 2pi/3

OpenStudy (alones.):

2pi/3 (or 2/3 pi) it's a ratio: pi radians = 180 degrees so 120 * (pi / 180) = pi * (120 / 180) = pi (2/3) or (2pi/3)

OpenStudy (alones.):

pi = 180 degrees. therefore, 120 degrees x (pi radians /180 degrees) = (2/3) pi radians because the degrees cancel out and you're left with radians.

OpenStudy (alones.):

Does this help?

OpenStudy (thcsbabygirl1224):

a little thank you

OpenStudy (alones.):

Yep! No problem!